OpenStudy (vexation):

JKLM is a rhombus. JL² + KM² = 4JK². M is the top left corner, L is in the top right, K is in the bottom right, and J is in the bottom left.

OpenStudy (mathmate):

Do you need a proof for the statement JL² + KM² = 4JK² ? In geometry, it is always a good idea to start with drawing a figure of the given situation. It is informative and inspiring.  Hints: If we inscribe the rhombus in a rectangle ABCD as shown above, we see that JL=BC and KM=AB. Show that JL^2+KM^2=AB^2+BC^2=AC^2. Finally, show the relation between AC and JK using similar triangles.